Au2S3 + 3 H2 = 2 Au + 3 H2S

How many formula units of gold (III) sulfide are required to produce 0.050 g of hydrosulfuric acid?



1
Expert's answer
2021-03-02T06:02:26-0500

n(H2S) = (0.050 g)/(34 g/mol) = 1.47*10-3 mol

n(Au2S3) = n(H2S)/3 = 4.90*10-4 mol

N(Au2S3) = n(Au2S3)*NA = (4.90*10-4 mol)*(6.02*1023 mol-1) = 2.95*1020 formula units